# ArtFormer: Controllable Generation of Diverse 3D Articulated Objects

### v1

@@ 6.3 Text Condition Generation Using GPT-4o
In some cases, GPT-4o may fail to produce a valid description (e.g., returning ”I’m sorry, I can’t assist with that.”), with a failure rate of $26.10\%$. In the final dataset for text-guided generation, such failed descriptions are excluded.
@@ 7.4 Image Guided Generation
Preliminary experiments of ArtFormer’s capability to generate articulated objects based on a single image, achieved by substituting the pretrained text encoder [47] with a pretrained image encoder [27], as discussed Fig. 6. Rendered images of articulated objects from PartNet-Mobility [61] using Blender, used as image conditions for generating articulated objects in the images with ArtFormer. The results are displayed in Fig. 15. In addition, we employ real-world photographs as the image condition to generate articulated objects. The results are illustrated in Fig. 16.

### v2

@@ 6.3 Text Condition Generation Using GPT-4o
In some cases, GPT-4o may fail to produce a valid description (e.g., returning "I’m sorry, I can’t assist with that."), with a failure rate of $26.10\%$. In the final dataset for text-guided generation, such failed descriptions are excluded.
@@ 7 Additional Experiments and Results
7.1 Novel Shape Generation
We conducted an experiment inspired by Diffusion-SDF to demonstrate that our shape prior, guided by an articulation transformer, can generate new geometry shapes that never appear in the dataset. We used our model to produce various objects and dissected them into sub-parts. Then, we calculated the Chamfer Distance between each sub-part and those in the training set and ranked them from nearest to farthest. The results, shown in Fig. 10, indicate that the sub-parts generated by our model are distinct from those in the training set, confirming the model’s ability to create novel geometry shapes.
@@ 7.4 Text Guided Generation
A quantitative evaluation is conducted to assess text-guided generation alignment, complementing previous human study results. Inspired by Park et al. [45], we adopt CLIP-R precision to measure the alignment between generated objects and instruction text. Snapshots are created for each generated object at an openness ratio of 0. The CLIP-R precision is then computed using these images and text instructions, leveraging the openai/clip-vit-large-patch14 model. The results for R=10 are presented in Tab. 3.
@@ 7.5 Image Guided Generation
Preliminary experiments of ArtFormer’s capability to generate articulated objects based on a single image, achieved by substituting the pretrained text encoder [51] with a pretrained image encoder [30], as discussed Sec. 4.4. Rendered images of articulated objects from PartNet-Mobility [65] using Blender, used as image conditions for generating articulated objects in the images with ArtFormer. The results are displayed in Fig. 15. In addition, we employ real-world photographs as the image condition to generate articulated objects. The results are illustrated in Fig. 16.
@@
Table 3: Text-Guided Object Alignment Results
CAGE NAP-128 NAP-768 Ours
CLIP-R@10$\uparrow$ 0.1429 0.1648 0.1319 0.2198